# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ 6 Conclusion
This paper contributes to out-of-distribution (OOD) statistical learning analysis by introducing bilevel optimization and upper bound minimization algorithms to actively optimize training data distributions for improved OOD generalization. Rooted in the theoretical analysis, these practical methods effectively learn training distributions over target sets, as demonstrated in a proof-of-concept using simple parametric families. While the current approach shows promise, it is limited by the use of low-dimensional parametric distribution classes and the global optimization step in the alternating algorithm. Future work will address these limitations by extending the methods to higher-dimensional and nonparametric probability distribution classes, such as transport maps and empirical measures. It is also worth refining these methods using derivative-informed Wasserstein gradient flows and tackling broader challenges, including cost–accuracy trade-offs for sample allocation and iteration complexity guarantees. This work provides a solid foundation for developing robust and scalable approaches to higher-dimensional OOD generalization problems in machine learning.
@@ Acknowledgments
N.G. and Y.Y. are partially supported by the U.S. National Science Foundation (NSF) through award DMS-2409855. N.H.N. acknowledges support from NSF award DMS-2402036. The authors further acknowledge the Amazon Web Services cloud computing grant provided by Cornell University’s Center for Data Science for Enterprise and Society.
Appendix A Additional mathematical background
@@
Example A.1 ($2$-Wasserstein metric between Gaussians).
Let $\mu=\mathcal{N}(m_{\mu},\mathcal{C}_{\mu})$ and $\nu=\mathcal{N}(m_{\nu},\mathcal{C}_{\nu})$ be Gaussian measures on a real separable Hilbert space $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$. Then [23] gives
$\mathsf{W}_{2}^{2}(\mu,\nu)=\lVert m_{\mu}-m_{\nu}\rVert^{2}+\operatorname{tr}\Big{(}\mathcal{C}_{\mu}+\mathcal{C}_{\nu}-2\bigl{(}\mathcal{C}_{\mu}^{1/2}\mathcal{C}_{\nu}\mathcal{C}_{\mu}^{1/2}\bigr{)}^{1/2}\Big{)}\,.$ (A.1)
The square root in (A.1) is the unique self-adjoint positive semidefinite (PSD) square root of a self-adjoint PSD linear operator. The trace of an operator $A\colon H\to H$ is given by $\operatorname{tr}(A)=\sum_{j\in\mathbb{N}}\langle e_{j},Ae_{j}\rangle$ for any orthonormal basis $\{e_{j}\}_{j\in\mathbb{N}}$ of $H$. If $A$ is PSD, then we can also write $\operatorname{tr}(A)=\lVert A^{1/2}\rVert_{\mathrm{HS}}^{2}$, where $B\mapsto\lVert B\rVert_{\mathrm{HS}}^{2}\coloneqq\sum_{j\in\mathbb{N}}\lVert Be_{j}\rVert^{2}$ is the squared Hilbert–Schmidt norm.
We frequently invoke (A.1) in this paper. When $H$ is infinite-dimensional, the Karhunen–Loève expansion (KLE) is often useful in calculations [51, Thm. 6.19, p. 533].
Lemma A.2 (KLE of Gaussian measure).
@@
$\operatorname{Law}\Biggl{(}m+\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}}\xi_{j}\phi_{j}\Biggr{)}=\mathcal{N}(m,\mathcal{C})\,.$ (A.2)
Generalizing Lem. A.2, there exist Karhunen–Loève expansions of non-Gaussian measures. In this case, the $\{\xi_{j}\}_{j\in\mathbb{N}}$ are non-Gaussian, have mean zero and variance one, and are only pairwise uncorrelated instead of independent and identically distributed (i.i.d.); still, the KLE “diagonalizes” the probability measure.
@@
Definition A.3 (spaces of probability measures).
@@
It is equipped with the $\sigma$-algebra $\mathcal{B}(\mathscr{P}(X))$ generated by the evaluation maps $e_{A}\colon\mathscr{P}(X)\to[0,1]$ given by $e_{A}(\mu)\coloneqq\mu(A)$ for all $A\in\mathcal{B}$. Similarly, the space of probability measures on $\mathscr{P}(X)$ is
@@
Def. A.3 can be related to the notion of random measure from Def. 2.2 as follows. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. Recall that a random probability measure on $X$ is a measurable map
@@
Example A.4 (empirical measure).
@@
Example A.5 (Dirichlet process).
@@
Example A.6 (random Gaussian measures).
In the infinite-dimensional state space setting, a canonical example is that of a random Gaussian measure defined through the KLE (A.2) of its mean and covariance. Let $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$ be infinite-dimensional, $\{\phi_{j}\}$ be an orthonormal basis of $H$, and $\{\lambda_{j}\}\subset\mathbb{R}_{\geq 0}$ and $\{\sigma_{ij}\}\subset\mathbb{R}_{\geq 0}$ be summable. Define the random measure $\mathcal{N}(m,\mathcal{C})$ by
$m\overset{\mathrm{Law}}{\coloneqq}\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}}\xi_{j}\phi_{j}\quad\text{and}\quad\mathcal{C}\coloneqq\mathcal{R}^{*}\mathcal{R},\quad\text{where}\quad\mathcal{R}\overset{\mathrm{Law}}{\coloneqq}\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\sqrt{\sigma_{ij}}\zeta_{ij}\phi_{i}\otimes\phi_{j}\,.$ (A.3)
The $\{\xi_{j}\}$ and $\{\zeta_{ij}\}$ are zero mean, unit variance, pairwise-uncorrelated real random variables. The outer product in (A.3) is defined by $(a\otimes b)c\coloneqq\langle b,c\rangle a$. Since $\mathcal{R}$ is a KLE in the space of Hilbert–Schmidt operators on $H$, it follows that $\mathcal{C}=\mathcal{R}^{*}\mathcal{R}$ is a valid self-adjoint PSD trace-class covariance operator with probability one. Since $\{\phi_{j}\}$ is fixed, the degrees of freedom in $\mathbb{Q}\coloneqq\operatorname{Law}(\mathcal{N}(m,\mathcal{C}))$ are the choice of eigenvalues $\{\lambda_{j}\}$ and $\{\sigma_{ij}\}$ and the distribution of the coefficients $\{\xi_{j}\}$ and $\{\zeta_{ij}\}$. Typically the coefficients are chosen i.i.d. from some common distribution such as the standard normal distribution $\mathcal{N}(0,1)$ or the uniform distribution on the interval $[-\sqrt{3},\sqrt{3}]$. The preceding construction of $\mathcal{N}(m,\mathcal{C})$ immediately extends to mixtures of Gaussian measures.
To conclude SM A, we connect back to the barycenters described in Rmk. 4.7.
Definition A.7 (Wasserstein barycenter).
@@
$\inf_{\mu\in\mathsf{P}}\operatorname{\mathbb{E}}_{\nu\sim\mathbb{Q}}\mathsf{W}_{p}^{p}(\mu,\nu)\,.$ (A.4)
@@
Example A.8 (Gaussian barycenter).
Let $\mathbb{Q}\coloneqq\operatorname{Law}(\mathcal{N}(m,\mathcal{C}))\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$, where $m\in\mathcal{U}$ and $\mathcal{C}\colon\mathcal{U}\to\mathcal{U}$ are the random mean vector and random covariance operator of a random Gaussian measure. Then the solution $\overline{\mu}$ to $\inf_{\mu\in\mathscr{P}_{2}(\mathcal{U})}\operatorname{\mathbb{E}}[\mathsf{W}_{2}^{2}(\mu,\mathcal{N}(m,\mathcal{C}))]$ is the Gaussian $\overline{\mu}=\mathcal{N}(\operatorname{\mathbb{E}}[m],\overline{\mathcal{C}})$ [4, Thm. 2.4], where $\overline{\mathcal{C}}$ is a non-zero PSD solution to the nonlinear equation
$\overline{\mathcal{C}}=\operatorname{\mathbb{E}}\biggl{[}\Bigl{(}\overline{\mathcal{C}}^{1/2}\mathcal{C}\overline{\mathcal{C}}^{1/2}\Bigr{)}^{1/2}\biggr{]}\,.$ (A.5)
One can employ fixed-point iterations to approximately solve (A.5) for $\overline{\mathcal{C}}$.
Appendix B Details for section 3Lipschitz theory for out-of-distribution generalization
This appendix provides the proof of Prop. 3.1 and expands on Ex. 3.3. We begin with the proof.
@@ Appendix B Details for section 3Lipschitz theory for out-of-distribution generalization
We begin by proving the first assertion, which is similar to the result and proof of [8, Lem. 7.2, p. 23]. Let $\phi(u)\coloneqq\lVert\mathcal{G}_{1}(u)-\mathcal{G}_{2}(u)\rVert$. For any $u$ and $u^{\prime}$, the triangle inequality yields
@@
Next, we return to Ex. 3.3 concerning Gaussian mixtures. The following lemma is used in this example; it is an independently interesting result.
Lemma B.1 (Lipschitz stability of mixtures in Wasserstein metric).
@@
In Lem. B.1, the Wasserstein distance $\mathsf{W}_{p}(\mathbb{Q},\mathbb{Q}^{\prime})$ uses $(\mathscr{P}_{p}(\mathcal{U}),\mathsf{W}_{p})$ as the ground metric space in Def. 2.1. Thus, it is a Wasserstein metric over the Wasserstein space.
Last, we remark on structural assumptions that allow us to further bound (3.5) in Ex. 3.3.
Remark B.2 (KLE mixing distributions).
Instate the setting and notation of Ex. 3.3. Furthermore, let $m\sim\Pi_{1}$ and $m^{\prime}\sim\Pi_{1}^{\prime}$ have centered (possibly non-Gaussian) KLEs
@@
In the preceding display, we further use the notation from (A.3) in Ex. A.6; thus, $\{\xi_{j}\}$ and $\{\xi_{j}^{\prime}\}$ have zero mean and unit variance, with laws in $\mathscr{P}_{2}(\mathbb{R})$. Let $\rho\in\Gamma(\eta,\eta^{\prime})$ be the $\mathsf{W}_{2}$-optimal coupling. Similar to [22, Lem. 4.1, pp. 13–14], let $\pi\in\Gamma(\Pi_{1},\Pi_{1}^{\prime})$ be a coupling with the property that $\pi=\operatorname{Law}(h,h^{\prime})$, where $h=\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}}z_{j}\phi_{j}$ and $h^{\prime}=\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}^{\prime}}z_{j}^{\prime}\phi_{j}$ in law and $(z_{j},z_{j}^{\prime})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho$. Then by (A.1) and the triangle inequality,
@@
This estimate shows that $\Pi_{1}$ and $\Pi_{1}^{\prime}$ are close in $\mathsf{W}_{2}$ if $\eta$ and $\eta^{\prime}$ are close in $\mathsf{W}_{2}$ and $\{\sqrt{\lambda_{j}}\}$ and $\{\sqrt{\lambda_{j}^{\prime}}\}$ are close in $\ell^{2}(\mathbb{N};\mathbb{R})$. We conclude this remark by upper bounding the term $\mathsf{W}_{2}^{2}(\Pi_{2},\Pi_{2}^{\prime})$ appearing (3.5). Following (A.3) from Ex. A.6, suppose that $\mathcal{C}=\mathcal{R}^{*}\mathcal{R}$ and $\mathcal{C}^{\prime}=(\mathcal{R}^{\prime})^{*}\mathcal{R}^{\prime}$, where
=$\mathcal{R}\overset{\mathrm{Law}}{=}\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\sqrt{\sigma_{ij}}\zeta_{ij}\phi_{i}\otimes\phi_{j},\,\mathcal{R}^{\prime}\overset{\mathrm{Law}}{=}\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\sqrt{\sigma_{ij}^{\prime}}\zeta_{ij}^{\prime}\phi_{i}\otimes\phi_{j},\,\zeta_{ij}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\varrho,\,\text{ and }\,\zeta_{ij}^{\prime}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\varrho^{\prime}$
are KLE expansions in the Hilbert space $\mathrm{HS}(\mathcal{U})$ of Hilbert–Schmidt operators from $\mathcal{U}$ into itself. Using (A.1) and [12, Lem. SM1.2, p. SM2], it holds that
@@
Appendix C Details for section 4Practical algorithms for training distribution design

### v3

@@ 6 Conclusion
This paper shows how optimizing training distributions using theoretically founded bilevel and alternating optimization algorithms produces more accurate and robust models for function approximation and PDE operator learning. Numerical experiments instate the methods with parametric distribution families and nonparametric particle-based discretizations of Wasserstein gradient flows. Though promising, the proposed approach is limited by distribution family expressivity, global optimization, and line search steps. Future work will address these limitations by incorporating transport maps and importance reweighting techniques. Understanding cost versus accuracy trade-offs for sample allocation and iteration complexity of the methodology is another important future challenge.
@@ Acknowledgments
N.G. and Y.Y. are partially supported by the U.S. National Science Foundation (NSF) through award DMS-2409855. N.H.N. acknowledges support from NSF award DMS-2402036 and from a Klarman Fellowship through Cornell University’s College of Arts & Sciences. The authors further acknowledge the Amazon Web Services cloud computing grant provided by Cornell University’s Center for Data Science for Enterprise and Society. The authors are also grateful to Kevin Miller for helpful comments on a previous version of the manuscript.
Supplementary Material for:
Learning where to learn: Training data distribution optimization for scientific machine learning
Appendix A Expanded overview of related work
Beyond the SciML literature reviewed in Sec. 1, our work also sits alongside research on learning under distribution shift [63]. This includes meta-learning, domain generalization and adaptation, and active learning and optimal experimental design (OED). We now explain these connections.
Meta-learning aims to learn models or priors that adapt rapidly across tasks drawn from a task distribution [44, 69]. While related in spirit, meta-learning typically treats the data-generating distribution per task as given; in contrast, we optimize the training distribution itself to improve average performance across deployment regimes. Both frameworks involve a two-level structure: an inner optimization that optimizes the model over a training dataset, and an outer optimization that adjusts a higher-level quantity beyond model parameters, guided by generalization performance. In optimization-based meta-learning methods, the outer loop adjusts the model initialization or update rules, allowing for rapid adaptation across new tasks. On the other hand, the outer loop in our work optimizes the training data distribution.
Domain generalization seeks representations that transfer across multiple source domains without access to target data [96, 25, 36, 30, 87, 98]. Instead, domain adaptation leverages (often unlabeled) target data to reduce source-target discrepancy [33]. Both paradigms assume a fixed training set and focus on achieving robustness during training. Common benchmarks for unsupervised domain adaptation only assume access to samples from a single target domain ($\nu^{\prime}$ in $\mathbb{Q}=\delta_{\nu^{\prime}}$) and do not have the ability to generate new labels [83]; this setting is orthogonal to our work. Robust domain generalization methods, such as distributionally robust optimization and invariant risk minimization [58, 78, 8, 51, 59, 101], modify objectives or constraints to ensure stability across domains given the training data. In contrast, we design the training distribution before learning: we retain standard training objectives but actively shape where data are drawn for improved sample complexity and robustness to distribution shift in regression problems. We do so with bilevel (or alternating) optimization in the space of probability measures, guided by theoretical bounds. In particular, the minimizing training data distribution depends not just on the chosen model architecture and the accuracy metric, but also on the target map itself.
The perspective presented in our paper naturally connects to active learning and OED. Active learning selects informative samples from a given data pool to improve label efficiency [26, 47, 71, 84, 100]; however, it typically operates reactively over a fixed unlabeled set during training. As a result, the support of the underlying distribution remains fixed as well. We instead study the a priori optimization of the sampling distribution itself. Like our work, there are active learning algorithms that are adaptive to labels [102]. However, this approach is pool-based, involves joint optimization instead of bilevel optimization, applies only to classification instead of regression, and adopts a worst-case instead of average-case perspective. On the other hand, OED explicitly reasons about where to make measurements, often through A- or D-optimality or Bayesian and sequential criteria [9, 34, 35, 45]. The classical OED problem emphasizes parameter inference and uncertainty reduction. In contrast, we target distributions that minimize the average OOD prediction error for function and operator learning, a design principle not commonly prioritized in OED. Nevertheless, our work shares similarities with modern measure-centric OED [43, 48], in which the design variable is a probability measure over the design space instead of a fixed number of sample points. In the nonlinear case, measure-based OED requires the formulation of a bilevel optimization problem [49]. The difference is that our inner loop involves the supervised training of a machine learning model, while for OED it typically involves solving a PDE-constrained inverse problem.
Other related work includes ridge leverage score and randomly pivoted Cholesky sampling for kernel regression [80, 18] as well as Bayesian optimization for sequential point-by-point dataset updates [85]. These techniques provide principled but typically label-independent sampling rules that do not depend on the target map to be learned. Instead, the present paper proposes adaptive, target-dependent, and supervised distribution design methods.
Appendix B Additional mathematical background
@@
Example B.1 ($2$-Wasserstein metric between Gaussians).
Let $\mu=\mathcal{N}(m_{\mu},\mathcal{C}_{\mu})$ and $\nu=\mathcal{N}(m_{\nu},\mathcal{C}_{\nu})$ be Gaussian measures on a real separable Hilbert space $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$. Then Gelbrich [39] gives
$\mathsf{W}_{2}^{2}(\mu,\nu)=\lVert m_{\mu}-m_{\nu}\rVert^{2}+\operatorname{tr}\Big(\mathcal{C}_{\mu}+\mathcal{C}_{\nu}-2\bigl(\mathcal{C}_{\mu}^{1/2}\mathcal{C}_{\nu}\mathcal{C}_{\mu}^{1/2}\bigr)^{1/2}\Big)\,.$ (B.1)
The square root in (B.1) is the unique self-adjoint positive semidefinite (PSD) square root of a self-adjoint PSD linear operator. The trace of an operator $A\colon H\to H$ is given by $\operatorname{tr}(A)=\sum_{j\in\mathbb{N}}\langle e_{j},Ae_{j}\rangle$ for any orthonormal basis $\{e_{j}\}_{j\in\mathbb{N}}$ of $H$. If $A$ is PSD, then we can also write $\operatorname{tr}(A)=\lVert A^{1/2}\rVert_{\mathrm{HS}}^{2}$, where $B\mapsto\lVert B\rVert_{\mathrm{HS}}^{2}\coloneqq\sum_{j\in\mathbb{N}}\lVert Be_{j}\rVert^{2}$ is the squared Hilbert–Schmidt norm.
We frequently invoke (B.1) in this paper. When $H$ is infinite-dimensional, the Karhunen–Loève expansion (KLE) is often useful in calculations [88, Thm. 6.19, p. 533].
Lemma B.2 (KLE of Gaussian measure).
@@
$\operatorname{Law}\Biggl(m+\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}}\xi_{j}\phi_{j}\Biggr)=\mathcal{N}(m,\mathcal{C})\,.$ (B.2)
Generalizing Lem. B.2, there exist Karhunen–Loève expansions of non-Gaussian measures. In this case, the $\{\xi_{j}\}_{j\in\mathbb{N}}$ are non-Gaussian, have mean zero and variance one, and are only pairwise uncorrelated instead of independent and identically distributed (i.i.d.); still, the KLE “diagonalizes” the probability measure.
@@
Definition B.3 (spaces of probability measures).
@@
It is equipped with the $\sigma$‑algebra $\mathcal{B}(\mathscr{P}(X))$ generated by the evaluation maps $e_{A}\colon\mathscr{P}(X)\to[0,1]$ given by $e_{A}(\mu)\coloneqq\mu(A)$ for all $A\in\mathcal{B}$. Similarly, the space of probability measures on $\mathscr{P}(X)$ is
@@
Def. B.3 can be related to the notion of random measure from Def. 2.2 as follows. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. Recall that a random probability measure on $X$ is a measurable map
@@
Example B.4 (empirical measure).
@@
Example B.5 (Dirichlet process).
@@
Example B.6 (random Gaussian measures).
In the infinite-dimensional state space setting, a canonical example is that of a random Gaussian measure defined through the KLE (B.2) of its mean and covariance. Let $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$ be infinite-dimensional, $\{\phi_{j}\}$ be an orthonormal basis of $H$, and $\{\lambda_{j}\}\subset\mathbb{R}_{\geq 0}$ and $\{\sigma_{ij}\}\subset\mathbb{R}_{\geq 0}$ be summable. Define the random measure $\mathcal{N}(m,\mathcal{C})$ by
$m\overset{\mathrm{Law}}{\coloneqq}\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}}\xi_{j}\phi_{j}\quad\text{and}\quad\mathcal{C}\coloneqq\mathcal{R}^{*}\mathcal{R},\quad\text{where}\quad\mathcal{R}\overset{\mathrm{Law}}{\coloneqq}\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\sqrt{\sigma_{ij}}\zeta_{ij}\phi_{i}\otimes\phi_{j}\,.$ (B.3)
The $\{\xi_{j}\}$ and $\{\zeta_{ij}\}$ are zero mean, unit variance, pairwise-uncorrelated real random variables. The outer product in (B.3) is defined by $(a\otimes b)c\coloneqq\langle b,c\rangle a$. Since $\mathcal{R}$ is a KLE in the space of Hilbert–Schmidt operators on $H$, it follows that $\mathcal{C}=\mathcal{R}^{*}\mathcal{R}$ is a valid self-adjoint PSD trace-class covariance operator with probability one. Since $\{\phi_{j}\}$ is fixed, the degrees of freedom in $\mathbb{Q}\coloneqq\operatorname{Law}(\mathcal{N}(m,\mathcal{C}))$ are the choice of eigenvalues $\{\lambda_{j}\}$ and $\{\sigma_{ij}\}$ and the distribution of the coefficients $\{\xi_{j}\}$ and $\{\zeta_{ij}\}$. Typically the coefficients are chosen i.i.d. from some common distribution such as the standard normal distribution $\mathcal{N}(0,1)$ or the uniform distribution on the interval $\sqrt{3},\sqrt{3}]$. The preceding construction of $\mathcal{N}(m,\mathcal{C})$ immediately extends to mixtures of Gaussian measures.
To conclude SM B, we connect back to the barycenters described in Rmk. 4.8.
Definition B.7 (Wasserstein barycenter).
@@
$\inf_{\mu\in\mathsf{P}}\operatorname{\mathbb{E}}_{\nu\sim\mathbb{Q}}\mathsf{W}_{p}^{p}(\mu,\nu)\,.$ (B.4)
@@
Example B.8 (Gaussian barycenter).
Let $\mathbb{Q}\coloneqq\operatorname{Law}(\mathcal{N}(m,\mathcal{C}))\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$, where $m\in\mathcal{U}$ and $\mathcal{C}\colon\mathcal{U}\to\mathcal{U}$ are the random mean vector and random covariance operator of a random Gaussian measure. Then the solution $\overline{\mu}$ to $\inf_{\mu\in\mathscr{P}_{2}(\mathcal{U})}\operatorname{\mathbb{E}}[\mathsf{W}_{2}^{2}(\mu,\mathcal{N}(m,\mathcal{C}))]$ is the Gaussian $\overline{\mu}=\mathcal{N}(\operatorname{\mathbb{E}}[m],\overline{\mathcal{C}})$ [6, Thm. 2.4], where $\overline{\mathcal{C}}$ is a non-zero PSD solution to the nonlinear equation
$\overline{\mathcal{C}}=\operatorname{\mathbb{E}}\biggl[\Bigl(\overline{\mathcal{C}}^{1/2}\mathcal{C}\overline{\mathcal{C}}^{1/2}\Bigr)^{1/2}\biggr]\,.$ (B.5)
One can replace the expectation with an empirical sum and employ fixed-point iterations to approximately solve (B.5) for $\overline{\mathcal{C}}$.
Appendix C Details for section 3Lipschitz theory for out-of-distribution error bounds
This appendix provides the proof of Prop. 3.1 and expands on Cor. 3.2 with Ex. C.1. We begin with the proof.
@@ Appendix C Details for section 3Lipschitz theory for out-of-distribution error bounds
We begin by proving the first assertion, which is similar to the result and proof of Benitez et al. [12, Lem. 7.2, p. 23]. Let $\phi(u)\coloneqq\lVert\mathcal{G}_{1}(u)-\mathcal{G}_{2}(u)\rVert$. For any $u$ and $u^{\prime}$, the triangle inequality yields
@@
Next, we apply Cor. 3.2 in the context of Gaussian mixtures.
Example C.1 (Gaussian mixtures).
Consider the random Gaussian measure $\nu^{\prime}=\mathcal{N}(m,\mathcal{C})\sim\mathbb{Q}\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$, where $(m,\mathcal{C})\sim\Pi$ and $\Pi$ is the mixing distribution over the mean (with marginal $\Pi_{1}$) and covariance (with marginal $\Pi_{2}$). Define the Gaussian mixture $\overline{\nu}\coloneqq\operatorname{\mathbb{E}}_{(m,\mathcal{C})\sim\Pi}\mathcal{N}(m,\mathcal{C})\in\mathscr{P}_{2}(\mathcal{U})$ via duality. Let $\overline{\mu}\coloneqq\operatorname{\mathbb{E}}_{(m^{\prime},\mathcal{C}^{\prime})\sim\Pi^{\prime}}\mathcal{N}(m^{\prime},\mathcal{C}^{\prime})$ be another Gaussian mixture with mixing distribution $\Pi^{\prime}$ (with marginals denoted by $\Pi_{1}^{\prime}$ and $\Pi_{2}^{\prime}$). By linearity of expectation and Fubini’s theorem, $\mathscr{E}_{\mathbb{Q}}(\mathcal{G})=\mathscr{E}_{\delta_{\overline{\nu}}}(\mathcal{G})$; that is, average-case accuracy with respect to $\mathbb{Q}$ is equivalent to accuracy with respect to the single test distribution $\overline{\nu}$; also see (D.2). This fact and Cor. 3.2 yield $\mathscr{E}_{\mathbb{Q}}(\mathcal{G})\leq\operatorname{\mathbb{E}}_{u\sim\overline{\mu}}\lVert\mathcal{G}^{\star}(u)-\mathcal{G}(u)\rVert^{2}+c(\mathcal{G}^{\star},\mathcal{G},\overline{\mu},\overline{\nu})\mathsf{W}_{2}(\overline{\mu},\overline{\nu})$. The factor $c(\mathcal{G}^{\star},\mathcal{G},\overline{\mu},\overline{\nu})$ can be further simplified using the closed-form expression for second moments of Gaussian measures. For the $\mathsf{W}_{2}$ factor, Lem. C.2 and coupling arguments imply that
$\mathscr{E}_{\mathbb{Q}}(\mathcal{G})\leq\operatorname{\mathbb{E}}_{u\sim\overline{\mu}}\big\lVert\mathcal{G}^{\star}(u)-\mathcal{G}(u)\big\rVert_{\mathcal{Y}}^{2}+c\bigl(\mathcal{G}^{\star},\mathcal{G},\overline{\mu},\overline{\nu}\bigr)\sqrt{\mathsf{W}_{2}^{2}(\Pi_{1},\Pi_{1}^{\prime})+\mathsf{W}_{2}^{2}(\Pi_{2},\Pi_{2}^{\prime})}\,.$ (C.1)
The two $\mathsf{W}_{2}$ distances between the marginals may be further bounded with (B.1) in SM B under additional structural assumptions on the mixing distributions; see Rmk. C.3. We often invoke this example in numerical studies with $\mathbb{Q}=\sum_{k=1}^{K}w_{k}\delta_{\nu_{k}^{\prime}}$ (where $K\in\mathbb{N}$, i.e., a finite mixture), weights $\sum_{k=1}^{K}w_{k}=1$, Gaussian test measures $\{\nu_{k}^{\prime}\}$, and degenerate training mixture $\overline{\mu}\coloneqq\mathcal{N}(m_{0},\mathcal{C}_{0})$.
The following lemma is used in Ex. C.1; it is an independently interesting result.
Lemma C.2 (Lipschitz stability of mixtures in Wasserstein metric).
@@
In Lem. C.2, the Wasserstein distance $\mathsf{W}_{p}(\mathbb{Q},\mathbb{Q}^{\prime})$ uses $(\mathscr{P}_{p}(\mathcal{U}),\mathsf{W}_{p})$ as the ground metric space in Def. 2.1. Thus, it is a Wasserstein metric over the Wasserstein space.
Last, we remark on structural assumptions that allow us to further bound (C.1) in Ex. C.1.
Remark C.3 (KLE mixing distributions).
Instate the setting and notation of Ex. C.1. Furthermore, let $m\sim\Pi_{1}$ and $m^{\prime}\sim\Pi_{1}^{\prime}$ have centered (possibly non-Gaussian) KLEs
@@
In the preceding display, we further use the notation from (B.3) in Ex. B.6; thus, $\{\xi_{j}\}$ and $\{\xi_{j}^{\prime}\}$ have zero mean and unit variance, with laws in $\mathscr{P}_{2}(\mathbb{R})$. Let $\rho\in\Gamma(\eta,\eta^{\prime})$ be the $\mathsf{W}_{2}$-optimal coupling. Similar to Garbuno-Inigo et al. [38, Lem. 4.1, pp. 13–14], let $\pi\in\Gamma(\Pi_{1},\Pi_{1}^{\prime})$ be a coupling with the property that $\pi=\operatorname{Law}(h,h^{\prime})$, where $h=\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}}z_{j}\phi_{j}$ and $h^{\prime}=\sum_{j\in\mathbb{N}}\sqrt{\lambda_{j}^{\prime}}z_{j}^{\prime}\phi_{j}$ in law and $(z_{j},z_{j}^{\prime})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho$. Then by (B.1) and the triangle inequality,
@@
This estimate shows that $\Pi_{1}$ and $\Pi_{1}^{\prime}$ are close in $\mathsf{W}_{2}$ if $\eta$ and $\eta^{\prime}$ are close in $\mathsf{W}_{2}$ and $\{\sqrt{\lambda_{j}}\}$ and $\{\sqrt{\lambda_{j}^{\prime}}\}$ are close in $\ell^{2}(\mathbb{N};\mathbb{R})$. We conclude this remark by upper bounding the term $\mathsf{W}_{2}^{2}(\Pi_{2},\Pi_{2}^{\prime})$ appearing (C.1). Following (B.3) from Ex. B.6, suppose that $\mathcal{C}=\mathcal{R}^{*}\mathcal{R}$ and $\mathcal{C}^{\prime}=(\mathcal{R}^{\prime})^{*}\mathcal{R}^{\prime}$, where
$\mathcal{R}\overset{\mathrm{Law}}{=}\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\sqrt{\sigma_{ij}}\zeta_{ij}\phi_{i}\otimes\phi_{j},\,\mathcal{R}^{\prime}\overset{\mathrm{Law}}{=}\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\sqrt{\sigma_{ij}^{\prime}}\zeta_{ij}^{\prime}\phi_{i}\otimes\phi_{j},\,\zeta_{ij}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\varrho,\,\text{ and }\,\zeta_{ij}^{\prime}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\varrho^{\prime}$
are KLE expansions in the Hilbert space $\mathrm{HS}(\mathcal{U})$ of Hilbert–Schmidt operators from $\mathcal{U}$ into itself. Using (B.1) and Cockayne and Duncan [20, Lem. SM1.2, p. SM2], it holds that
@@
Appendix D Details for section 4Practical algorithms for training distribution design